OpenStudy (anonymous):

may any one help me plz in this exercise A 50-mg dose of a drug is administered by rapid injection to a patient. The half-life of the drug is 9 hours. a. find a model for the amount of drug in the bloodstream t hours after the drug is administered.

OpenStudy (rogue):

\[F(t) = A e^{-kt}\]\[k = \frac {\ln 2}{t_{1/2}} = \frac {\ln 2}{9} \approx 0.07710635\]\[A = Initial Amount = 50mg\]\[F(t) = 50 e^{-0.0771t}\]